A line passes through (1,-5) and (-3,7) write an equation for the line in point - slope form
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

x1 y1 x2 y2
1 -5 -3 7

slope m = (y2-y1)/(x2-x1)
( 7 - -5 )/( -3 - 1 )
( 12 / -4 )
m= - 3

Plug value of the slope and point ( -1 , 8 ) in
Y = m x + b
8.00 = -3 + b
b= 11
So the equation will be
Y = -3 x + 11
m.ananth@hotmail.ca

Answer by Alan3354(69443)

A line passes through (1,-5) and (-3,7) write an equation for the line in point - slope form
-------------
m = diffy/diffx = 12/-4
m = -3
---------
y = mx + b
Use either point
-5 = -3 + b
b = -2
---------
y = -3x - 2